OpenStudy (anonymous):

Find the present value of $10,000 if interest is paid at a rate of 5% per year, compounded weekly, for 5 years. (Round your answer to the nearest cent.)

OpenStudy (kropot72):

\[\large A=10000(1+\frac{0.05}{52})^{(5\times52)}\]

OpenStudy (anonymous):

this os for my online homework and it keeps getting mark wrong when i put my answer of 12838.71 in

OpenStudy (kropot72):

Perhaps the question is not well worded. Is it looking for the original principal, when the amount after 5 years is $10,000. \[\large 10000=P(1+\frac{0.05}{52})^{(5\times52)}\]

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

I do solve this if it is looking for the original principle

OpenStudy (kropot72):

If the revised interpretation is correct, the "present value" would be $(10,000/1.28387), this being the amount needed now to give $10,000 in 5 years time.